My mum was 27 when I was born. 8 years ago she was twice as old as I shall be in 5 years time. How old am I now?
Answer by ankor@dixie-net.com(22740)

My mum was 27 when I was born.
8 years ago she was twice as old as I shall be in 5 years time.
How old am I now?
:
m = mother's present age
c = child's present age
:
"My mum was 27 when I was born.:
From this statement:
m = c + 27
:
"8 years ago she was twice as old as I shall be in 5 years time."
m - 8 = 2(c + 5)
m - 8 = 2c + 10
m = 2c + 10 + 8
m = 2c + 18
;
How old am I now?
Replace m with (c+27) in the above equation
c + 27 = 2c + 18
27 - 18 = 2c - c
c = 9 yrs
:
;
Check solution: m = 9 + 27 = 36
"8 years ago she was twice as old as I shall be in 5 years time."
36 - 8 = 2(9 + 5)
28 = 2(14)